OpenStudy (mikeymaximum):

There are 7 oak trees in Aidan's yard. The other 2 trees in his yard are maple trees. What fraction of the trees in Aidan's yard are oak trees? A. 9/7 B. 2/7 C. 7/9 D. 2/9

OpenStudy (mikeymaximum):

medal and fan

OpenStudy (mikeymaximum):

@pooja195

OpenStudy (k_lynn):

You're going to need to find out how many trees there are all together and then create a fraction. So what's 7+2? After you figure that out, you will but the number of oak trees as the numerator and the overall amount of trees as the denominator. \[\frac{ Number~of~oak~trees }{ Number~of~all~trees }\]

OpenStudy (anonymous):

total trees = 7+2 = 9 Oak trees = 7 fraction = oak trees/ total trees = 7/ 9

OpenStudy (k_lynn):

Yep, you got it @akitav